### v1

# Growth of Gravitational Wave Spectrum from Sound Waves in a Universe with Generic Expansion Rate

Abstract
We derived here the factor $\Upsilon$, which quantifies how the gravitational wave spectrum generated by sound waves in the radiation sector grows over time, in a universe with a generic expanding rate set by another dominant energy content. When the dominant energy density satisfies $\rho\propto a^{-3(1+w)}$, we found that $\Upsilon$ has a compact analytical expression: $\Upsilon=\frac{2[1-y^{3(w-1)/2}]}{3(1-w)}$, where $y=a(t)/a(t_{s})$ which is the ratio of the scale factor at a later time $t$ to that at $t_{s}$ when gravitational wave production from sound waves starts. This generic result reduces to that derived previously for radiation-dominated and matter-dominated cases, thus generalizing previous formulas to more general cosmological contexts and providing more accurate results. The derivation relies solely on a stationary source, implying that this generic result of $\Upsilon$ serves as an universal factor in describing the growth of the gravitational wave production and can appear beyond cosmological phase transitions.
preprint: APS/123-QED
I Introduction
The stochastic background of gravitational waves from cosmological first-order phase transitions is one of the promising frontiers for discovering physics beyond the Standard Model. LIGO data has been used to search for gravitational waves originating from new physics at energy scales of $\mathcal{O}(10^{3}\sim 10^{6})$ TeV Romero et al. (2021); Badger et al. (2023), far exceeding the energy scale accessible by the current large colliders. Additionally, pulsar timing array experiments reported evidences for a stochastic gravitational wave background, indicating QCD phase transitions or strongly supercooled phase transitions as possible sources Xue et al. (2021); Arzoumanian et al. (2021); Agazie et al. (2023); Antoniadis et al. (2023, 2024); Xiao et al. (2023); Han et al. (2024); Ahmadvand et al. (2023); Zheng et al. (2024); Bian et al. (2021, 2024). Space-based gravitational waves detectors, such as LISA Amaro-Seoane et al. (2017), Taiji Hu and Wu (2017); Ruan et al. (2020) and Tianqin Luo et al. (2016); Mei et al. (2021); Luo et al. (2020), are currently in preparation and expected to come into play in the coming decades.
To ascertain whether the detected gravitational waves indeed originated from the early universe phase transitions, a profound understanding of the relevant processes is essential. It is widely believed that phase transition gravitational waves can primarily be generated through three mechanisms: bubble wall collisions Kosowsky and Turner (1993); Kosowsky et al. (1992); Jinno and Takimoto (2017); Zhao et al. (2023); Di et al. (2021); Li et al. (2023), sound waves Hindmarsh et al. (2014, 2015), and magnetohydrodynamic (MHD) turbulence Caprini et al. (2009); Kahniashvili et al. (2008, 2010); Caprini et al. (2016); Roper Pol et al. (2020). Gravitational waves produced by bubble walls can be well-described by the envelope approximation, and an analytical expression has been obtained Jinno and Takimoto (2017). In thermal plasma phase transitions, gravitational waves induced by sound waves are dominant and have been extensively studied, while those from MHD turbulence are sub-dominant and related studies remain highly uncertain. Although the behavior of gravitational waves induced by sound waves can be reasonably described by the sound shell model, simulation results suggest the presence of additional mechanisms affecting the final spectrum Cutting et al. (2020). This indicates a need for further investigation into this process. Recently, Ref. Cai et al. (2023) focused on the forced propagating contribution from the initial collision stage of sound shells and achieved better agreement with simulations compared to the sound shell model. Ref. Ellis et al. (2020) compared several models and discovered that the parameter regions with high signal-to-noise ratios (SNR) often coincide with short duration of sound waves, potentially reducing the observable gravitational wave spectrum. In our previous work Guo et al. (2021), we more carefully calculated the gravitational wave spectrum of the sound shell model, and was able to extract an overall factor connected with the duration of sound waves, called $\Upsilon$. Ref. Roper Pol et al. (2024) pointed out that some assumptions in the sound shell model lead to incorrect infrared behavior and extended the $\Upsilon$ factor to the infrared regime. However, most of the above calculations are in the radiation-dominated universe and have not been verified in other universe scenarios. If the equation of state of the universe deviates from the ultra-relativistic case during a phase transition, the gravitational wave spectrum may be modified Giombi et al. (2024). Therefore, it is both necessary and interesting to extend these results to a more general cosmological period to reproduce the potential gravitational wave generation processes as accurately as possible.
The aim of this work is to follow our previous work and try to extend the $\Upsilon$ factor from the radiation-dominated universe and the matter-dominated universe to an arbitrary single-component dominated universe. The structure of the paper is as follows: in section II, we present the theoretical framework for the sound shell model, followed by our extensions in section III. In section IV, we provide the resulting gravitational wave spectrum, and finally, the conclusions are discussed in section V.
II Theoretical Framework and Sound Shell Model
We first consider the gravitational waves in the FLRW universe, described by the metric
${\rm d}s^{2}=-{\rm d}t^{2}+a(t)^{2}\left[\delta_{ij}+h_{ij}\right]{\rm d}\mathbf{x}^{2},$ (1)
where $h_{ij}$ is the transverse traceless part of the tensor-mode perturbations. Gravitational waves are sourced by the transverse traceless part of the perturbed stress-energy tensor of the matter content $\pi_{ij}^{T}(t,\mathbf{x})$. Those two quantities are connected by the Einstein equation, which could be written in Fourier space as
$h^{\prime\prime}_{ij}(t,\mathbf{q})+2\frac{a^{\prime}}{a}h_{ij}^{\prime}(t,\mathbf{q})+q^{2}h_{ij}(t,\mathbf{q})=16\pi Ga^{2}\pi_{ij}^{T}(t,\mathbf{q}),$ (2)
where $\prime$ represents a derivative with respect to conformal time $\eta$ to distinguish it from the one with respect to coordinate time, denoted by a dot. The gravitational wave energy density is defined as
$\rho_{\rm GW}=\frac{\left<\dot{h}_{ij}(t,\mathbf{x})\dot{h}_{ij}(t,\mathbf{x})\right>}{32\pi G},$ (3)
[…]
$\dot{h}_{ij}(t,\mathbf{q}_{1})\dot{h}_{ij}(t,\mathbf{q}_{2})\right>=(2\pi)^{2}\delta^{3}(\mathbf{q}_{1}+\mathbf{q}_{2})P_{\dot{h}}(q_{1},t).$ (4)
Then the gravitational wave energy density can be reorganized as
$\rho_{\rm GW}=\frac{1}{64\pi^{3}G}\int{\rm d}q\ q^{2}P_{\dot{h}}(q,t),$ (5)
and the gravitational wave energy density spectrum follows
$\frac{{\rm d}\rho_{\rm GW}(t)}{{\rm d}{\rm ln}q}=\frac{q^{3}P_{\dot{h}}(t,q)}{64G\pi^{3}}.$ (6)
In practice, the energy density is often used in the form of dimensionless fraction $\Omega_{\rm GW}(t)=\rho_{\rm GW}(t)/\rho_{\rm c}(t)$, where $\rho_{\rm c}$ is the critical energy density at time $t$. So the dimensionless version of the spectrum is
$\mathcal{P}_{\rm GW}\equiv\frac{{\rm d}\Omega_{\rm GW}(t)}{{\rm d}{\rm ln}q}=\frac{q^{3}P_{\dot{h}}(t,q)}{24\pi^{2}H^{2}}=\frac{q^{3}P_{h^{\prime}}(t,q)}{24\pi^{2}H^{2}a^{2}}.$ (7)
The gravitational wave equation Eq.(2) could be solved by the Green’s function with the following boundary conditions
$G(\tilde{\eta}\leq\tilde{\eta}_{0})=0,$ (8)
$\left.\frac{\partial G(\tilde{\eta},\tilde{\eta}_{0})}{\partial\tilde{\eta}}\right|_{\tilde{\eta}=\tilde{\eta}_{0}^{+}}=1,$ (9)
where $\tilde{\eta}=q\eta$ is a dimensionless quantity and the subscript zero indicates the time when the phase transition begins. The solution is then given by
$h_{ij}(t,\mathbf{q})=16\pi G\int_{\tilde{\eta}_{0}}^{\tilde{\eta}}{\rm d}\tilde{\eta}^{\prime}G(\tilde{\eta},\tilde{\eta}^{\prime})\frac{a^{2}(\tilde{\eta}^{\prime})\pi_{ij}^{T}(\tilde{\eta}^{\prime},q)}{q^{2}}.$ (10)
Inserting the above solution into the left hand of Eq.(4), we could obtain
$P_{h^{\prime}}(q_{1},\eta)$ $=(16\pi G)^{2}\int_{\tilde{\eta}_{0}}^{\tilde{\eta}}{\rm d}\tilde{\eta}_{1}\int_{\tilde{\eta}_{0}}^{\tilde{\eta}}{\rm d}\tilde{\eta}_{2}\left\{\frac{\partial G(\tilde{\eta},\tilde{\eta}_{1})}{\partial\tilde{\eta}}\frac{\partial G(\tilde{\eta},\tilde{\eta}_{2})}{\partial\tilde{\eta}}\right.$
×$\left.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{a^{2}(\eta_{1})a^{2}(\eta_{2})}{q^{2}}\Pi^{2}(q_{1},\eta_{1},\eta_{2})\right\},$ (11)
where we have used the UETC of $\pi_{ij}^{T}$
$\pi_{ij}^{T}(\eta_{1},\mathbf{q}_{1})\pi_{ij}^{T}(\eta_{2},\mathbf{q}_{2})\right>=(2\pi)^{3}\delta^{3}(\mathbf{q}_{1}+\mathbf{q}_{2})\Pi^{2}(q_{1},\eta_{1},\eta_{2}).$ (12)
The transverse-traceless part of the energy-momentum tensor, $\pi_{ij}^{T}$, is given by
$\pi_{ij}^{T}=\Lambda_{ij,kl}\left(\bar{p}+\bar{e}\right)v^{k}v^{l},$ (13)
where $\Lambda_{ij,kl}$ is the transverse-traceless projection operator, $\bar{p}$ and $\bar{e}$ are the averaged pressure and energy density. Then we could express the $\Pi^{2}$ as the four point correlation function of the fluid velocity. With the following Fourier transformation of the Gaussian random field of the fluid velocity $\tilde{v}^{i}_{\mathbf{q}}(\eta)$,
$\tilde{v}_{\mathbf{q}}^{i}(\eta)=\int{\rm d}^{3}x~{}v^{i}(\mathbf{x},\eta)e^{-i\mathbf{q}\cdot\mathbf{x}},\$ (14)
and according to the Wick theorem, $\Pi^{2}$ could be simplified as follows
$\Pi^{2}(k,\eta_{1}-\eta_{2})=\left(\bar{e}+\bar{p}\right)^{2}\int$ $\frac{{\rm d}^{3}q}{(2\pi)^{3}}\bigg{\{}G(q,\eta_{1},\eta_{2})G(\tilde{q},\eta_{1},\eta_{2})$
×$\left.\times\frac{q^{2}}{\tilde{q}^{2}}(1-\mu^{2})^{2}\right\},$ (15)
where $\tilde{q}=\left|\mathbf{q}-\mathbf{k}\right|$ and $\mu=\hat{\mathbf{q}}\cdot\hat{\mathbf{k}}$. In the derivation of the above expression, we have used the UETC of the Fourier coefficients of fluid velocity
$\tilde{v}^{i}_{\mathbf{q}}(\eta_{1})\tilde{v}^{j*}_{\mathbf{k}}(\eta_{2})\right>=\delta^{3}(\mathbf{q}-\mathbf{k})\hat{q}^{i}\hat{k}^{j}G(q,\eta_{1},\eta_{2}),$ (16)
and the property of the projection operator
$\Lambda_{ij,kl}\hat{q}^{i}\hat{\tilde{q}}^{j}\hat{q}^{k}\hat{\tilde{q}}^{l}=\frac{q^{2}}{\tilde{q}}\frac{(1-\mu^{2})^{2}}{2}.$ (17)
Note that the source of the gravitational wave here is approximately stationary. This implies that $\Pi^{2}$ only depends on $\eta_{1}-\eta_{2}$, and leads to a factorized form for the spectrum, corresponding to the $\Upsilon$ factor Guo et al. (2021).
In order to explicitly display the dependence on the scale factor, we define the rescaled and dimensionless quantity $\tilde{\Pi}$ as Guo et al. (2021)
$\Pi^{2}(q,\eta_{1},\eta_{2})=\frac{a_{s}^{8}}{a^{4}(\eta_{1})a^{4}(\eta_{2})}\left[\left(\bar{\tilde{e}}+\bar{\tilde{p}}\right)\bar{U}_{f}^{2}\right]^{2}L_{f}^{3}\tilde{\Pi}^{2}(qL_{f},q\eta_{1},q\eta_{2}),$ (18)
where $a_{s}$ is the scale factor when the source begins active, $\bar{\tilde{e}}$ and $\bar{\tilde{p}}$ are the rescaled averaged energy density and pressure, $\bar{U}_{f}$ is the magnitude of the fluid velocity and $L_{f}$ is the characteristic length during the phase transition, which is usually chosen as the mean bubble separation $R_{*}$.
In a first-order phase transition, the expanding vacuum bubble wall exerts an influence on the surrounding fluid. After bubble collisions, we expect the fluid to propagate like sound waves and become the dominant source for gravitational waves production. The calculation of $G(q,\eta_{1},\eta_{2})$ requires obtaining the specific form of $\tilde{v}_{\mathbf{q}}^{i}$ after bubble collision, in other words, we need to establish a relationship between the fluid velocity before the bubble collision ($\eta<\eta_{c}$) and that after the bubble collision ($\eta>\eta_{c}$). Because the vacuum bubbles are nucleated randomly, it is easier to look at only one bubble first, for example the $n$-th bubble. We could formally decompose this velocity field before collision as
$v^{i(n)}(\eta\leq\eta_{c}^{(n)},\mathbf{x})$ $=\frac{1}{2}\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}\left[\tilde{v}^{i(n)}_{\mathbf{q}}(\eta)e^{i\mathbf{q}\cdot\mathbf{x}}\right.$
$\left.~{}~{}~{}~{}~{}~{}~{}~{}+\tilde{v}^{i(n)*}_{\mathbf{q}}(\eta)e^{-i\mathbf{q}\cdot\mathbf{x}}\right].$ (19)
After bubble collision, the sound waves freely propagate and admit the following decomposition
$v^{i(n)}(\eta\geq\eta_{c}^{(n)},\mathbf{x})$ $=\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}\left[v_{\mathbf{q}}^{i(n)}e^{-i\omega\eta+i\mathbf{q}\cdot\mathbf{x}}\right.$
$\left.~{}~{}~{}~{}~{}~{}~{}~{}+v_{\mathbf{q}}^{i(n)*}e^{i\omega\eta-i\mathbf{q}\cdot\mathbf{x}}\right],$ (20)
where $\omega=c_{s}q$. To match the above two different processes, it is natural to evaluate those equations at collision time $\eta_{c}$ simultaneously and we will find that
$v_{\mathbf{q}}^{i(n)}=\frac{1}{2}\left[\tilde{v}_{\mathbf{q}}^{i(n)}(\eta_{c}^{(n)})+\frac{i}{\omega}\tilde{v}_{\mathbf{q}}^{i(n)\prime}(\eta_{c}^{(n)})\right]e^{i\omega\eta_{c}^{(n)}}.$ (21)
By the Euler equation of the fluid
$\tilde{v}_{\mathbf{q}}^{j\prime}+ic_{s}^{2}q^{j}\tilde{\lambda}_{\mathbf{q}}=0,$ (22)
where $\tilde{\lambda}_{\mathbf{q}}$ is the Fourier coefficients of the energy fluctuation $\lambda=\left[\tilde{e}(x)-\bar{\tilde{e}}\right]/\bar{\tilde{w}}$, we could write the Fourier coefficients of the sound waves as
$v_{\mathbf{q}}^{i(n)}=\frac{1}{2}\left[\tilde{v}_{\mathbf{q}}^{i(n)}(\eta_{c}^{(n)})+c_{s}\hat{q}^{i(n)}\tilde{\lambda}_{\mathbf{q}}(\eta_{c}^{(n)})\right]e^{i\omega\eta_{c}^{(n)}}.$ (23)
According to the assumption of the sound shell model, the total velocity and energy density field are the superposition of the corresponding fields produced by the $N_{b}$ expanding bubblesHindmarsh (2018); Hindmarsh and Hijazi (2019)
$v_{\mathbf{q}}^{i}=\sum_{n}^{N_{b}}v_{\mathbf{q}}^{i(n)}.$ (24)
The velocity field configuration produced by a single bubble before bubble collision can be obtained using the standard energy budget method Espinosa et al. (2010). It is straightforward to relate this to its corresponding Fourier coefficients
$\tilde{v}_{\mathbf{q}}^{j(n)}(\eta_{c}^{(n)})$ $=ie^{-i\mathbf{q}\cdot\mathbf{x}^{(n)}}\left[T^{(n)}\right]^{3}\hat{z}^{(n)j}f^{\prime}(z^{(n)}),$ (25)
$\tilde{\lambda}_{\mathbf{q}}^{(n)}(\eta_{c}^{(n)})$ $=e^{-i\mathbf{q}\cdot\mathbf{x}^{(n)}}\left[T^{(n)}\right]^{3}l(z^{(n)}),$ (26)
where $T^{(n)}\equiv\eta_{c}^{(n)}-\eta_{\rm nucleate}^{(n)}$ represents the lifetime of the $n$-th bubble, and $z^{n}\equiv\mathbf{q}T^{(n)}$. The specific form of $f(z)$ and $l(z)$ can be seen in Ref.Guo et al. (2021); Hindmarsh and Hijazi (2019). Inserting those into Eq.(23), we finally get
$v_{\mathbf{q}}^{i(n)}=i\hat{z}^{(n)i}\left[T^{(n)}\right]^{3}e^{i\omega\eta_{c}^{(n)}-i\mathbf{q}\cdot\mathbf{x}^{(n)}}A(z^{n}),$ (27)
where $A(z)=\left[f^{\prime(z)}-ic_{s}l(z)\right]/2$.
Once we have obtained $v^{i(n)}_{\mathbf{q}}$, we could directly compute $\tilde{v}_{\mathbf{q}}^{i}(\eta\geq\eta_{c}^{(n)})$ by its definition. After doing the ensemble average over nucleation sites of all bubbles, the velocity power spectrum $G(q,\eta_{1},\eta_{2})$ is
$G(q,\eta_{1},\eta_{2})=2P_{v}(q){\rm cos}\left[\omega(\eta_{1}-\eta_{2})\right],$ (28)
where $P_{v}$ contains the integration of $A(z)$, and its specific form can be also found in Ref.Guo et al. (2021); Hindmarsh and Hijazi (2019). It is worth noting that Ref.Roper Pol et al. (2024) pointed out that there is an extra term in Eq.(28), which is important to reproduce the causality condition in the infrared region.
III $\Upsilon$ factor in singlet component universe
Now we have gathered all the components except for the Green’s function. Since we aim to obtain the $\Upsilon$ factor for arbitrary single-component background universes, we need to extend the calculation of the Green’s function into a more general case.
By the continuity equation for a perfect fluid
$\dot{\rho}+\frac{3\dot{a}}{a}(\rho+P)=0,$ (29)
and the constant equation of state between pressure and energy density $P=w\rho$, the scale factor could relate to the energy density as
$\rho\propto a^{-3(1+w)}.$ (30)
Then the different $w$ corresponds to the different single-component matter which dominates in the universe. We treat $w$ as a fixed input constant in this paper, ignoring variations in this parameter. This assumption is valid if the timescale of the gravitational wave production is relatively short.
Inserting Eq.(30) into the Friedmann equation of flat universe
$\frac{{\rm d}a}{{\rm d}\eta}=a^{2}\sqrt{\frac{8\pi G}{3}\rho}=a^{2}\sqrt{H_{s}^{2}\left(\frac{a}{a_{s}}\right)^{-3(1+w)}},$ (31)
and defining $y(\eta)=a(\eta)/a_{s}$, following our previous derivation Guo et al. (2021), we could obtain
$\frac{{\rm d}y}{{\rm d}\eta}=a_{s}H_{s}y^{(1-3w)/2},$ (32)
where $H_{s}$ is the Hubble constant when the source begins active. With the initial condition $y(\eta_{s})=1$, the solution of the above equation is
$y^{\frac{1+3w}{2}}=\frac{3w+1}{2}(\eta-\eta_{s})H_{s}a_{s}+1.$ (33)
Changing the variable from $\eta$ to $y$, the gravitational wave equation Eq.(2) could be rewritten as
$\frac{{\rm d}^{2}h_{ij}}{{\rm d}y^{2}}+\frac{{\rm d}h_{ij}}{{\rm d}y}\left(\frac{5-3w}{2y}\right)+\frac{q^{2}y^{3w-1}}{(H_{s}a_{s})^{2}}h_{ij}=\frac{16\pi Gy^{3w-1}a^{2}}{(H_{s}a_{s})^{2}}\pi_{ij}^{T}.$ (34)
Then denoting $q^{2}/(H_{s}a_{s})^{2}=\tilde{q}^{2}$ and $\tilde{y}=y\tilde{q}^{n}$, the equation could be simplified further
$\tilde{y}^{1-3w}\frac{{\rm d}^{2}h_{ij}}{{\rm d}\tilde{y}^{2}}+\tilde{y}^{-3w}\frac{{\rm d}h_{ij}}{{\rm d}\tilde{y}}(\frac{5-3w}{2})+h_{ij}=\frac{16\pi Ga^{2}}{q^{2}}\pi_{ij}^{T},$ (35)
by setting $2n=2-n(3w-1)$. The corresponding Green’s function is the homogeneous solution of this equation and could be expressed as Bowman (2012)
$G(\tilde{y},\tilde{y}_{0})$ $=\tilde{y}^{\frac{3w-3}{4}}\left[C_{1}J_{\nu}\left(\frac{2}{3w+1}\tilde{y}^{\frac{3w+1}{2}}\right)\right.$
$\left.+C_{2}J_{-\nu}\left(\frac{2}{3w+1}\tilde{y}^{\frac{3w+1}{2}}\right)\right],$ (36)
where $J_{\nu}(x)$ is the Bessel function of the first kind and the $\nu=(3w-3)/(6w+2)$ should not be an integer. If $\nu$ is an integer, the $J_{-\nu}$ in the above expression should be replaced with the Bessel function of the second kind to satisfy the condition of linear independence Bowman (2012). In this paper, we confine ourselves to values of $w$ that satisfy the strong-energy condition, i.e., $01+3w>$ 0, which implies that almost all values of $\nu$ are not integers. The constants $C_{1}$ and $C_{2}$ can be determined by the following initial conditions
$G(\tilde{y}\leq\tilde{y}_{0})=0,$ (37)
$\left.\frac{\partial G(\tilde{y},\tilde{y}_{0})}{\partial\tilde{y}}\right|_{\tilde{y}=\tilde{y}_{0^{+}}}=\tilde{y}_{0}^{3w+1},$ (38)
with the following results:
$C_{1}$ $=-\frac{b}{b_{1}a-ba_{1}}\tilde{y}_{0}^{\frac{3w+1}{4}},$ (39)
$C_{2}$ $=\frac{a}{b_{1}a-ba_{1}}\tilde{y}_{0}^{\frac{3w+1}{4}},$ (40)
where $a=J_{\nu}\left(\frac{2}{3w+1}\tilde{y}_{0}^{\frac{3w+1}{2}}\right)$, $b=J_{-\nu}\left(\frac{2}{3w+1}\tilde{y}_{0}^{\frac{3w+1}{2}}\right)$ and $a_{1}$ ($b_{1}$) is the corresponding derivative with respect to its variable, ${\rm d}J_{\nu}(z)/{\rm d}z$ (${\rm d}J_{-\nu}(z)/{\rm d}z$).
With the Green’s function, the dimensionless gravitational wave spectrum $\mathcal{P}_{\rm GW}$ is given by
$\mathcal{P}_{\mathrm{GW}}\left(y,qR_{*}\right)=\frac{\left[16\pi G\left(\bar{\tilde{e}}+\bar{\tilde{p}}\right)\bar{U}_{f}^{2}\right]^{2}\left(qR_{*}\right)^{3}}{24\pi^{2}H^{2}H_{s}^{2}y^{3w+1}}$
×$\times\int_{y_{s}}^{y}{\rm d}y_{1}\int_{y_{s}}^{y}{\rm d}y_{2}\left\{\left(\frac{q}{a_{s}H_{s}}\right)^{4n-2}\frac{\partial G\left(\tilde{y},\tilde{y}_{1}\right)}{\partial\tilde{y}}\frac{\partial G\left(\tilde{y},\tilde{y}_{2}\right)}{\partial\tilde{y}}\right.$
×$\left.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{1}{y_{1}^{2}y_{2}^{2}}\frac{\tilde{\Pi}^{2}\left(qR_{*},q|\eta_{1}-\eta_{2}|\right)}{\tilde{q}^{2}}\right\}.$ (41)
Because the source is largely stationary, the dependence of $\mathcal{P}_{\rm GW}$ on the conformal time is only through the difference
$q|\eta_{1}-\eta_{2}|=\frac{2qR_{*}}{R_{*}a_{s}H_{s}(3w+1)}\left|y_{1}^{\frac{3w+1}{2}}-y_{2}^{\frac{3w+1}{2}}\right|.$ (42)
This results in a factorized form of the spectrum $\mathcal{P}_{\rm GW}$ for most frequencies in the range near the peak Guo et al. (2021). This factor, called $\Upsilon$ in previous studies, can be obtained in the following way.
Firstly we change the integral variables from $y_{1}$ and $y_{2}$ to $y_{+}$ and $y_{-}$
$y_{-}$ $=\frac{2}{3w+1}(y_{1}^{\frac{3w+1}{2}}-y_{2}^{\frac{3w+1}{2}}),$ (43)
$y_{+}$ $=\frac{1}{3w+1}(y_{1}^{\frac{3w+1}{2}}+y_{2}^{\frac{3w+1}{2}}).$ (44)
The corresponding Jacobian determinant is $(y_{1}y_{2})^{\frac{1-3w}{2}}$ and the new limits of integration are
$0\leq y_{-}\leq(y^{\frac{3w+1}{2}}-1)\frac{2}{3w+1},$ (45)
$\frac{2}{3w+1}+\frac{y_{-}}{2}\leq y_{+}\leq-\frac{y_{-}}{2}+\frac{2}{3w+1}y^{\frac{3w+1}{2}},$ (46)
and
$(1-y^{\frac{3w+1}{2}})\frac{2}{3w+1}\leq y_{-}\leq 0,$ (47)
$-\frac{y_{-}}{2}+\frac{2}{3w+1}\leq y_{+}\leq\frac{y_{-}}{2}+y^{\frac{3w+1}{2}}\frac{2}{3w+1}.$ (48)
To explicitly show the $a^{-4}$ behavior of the gravitational radiation, we also denote
$\frac{\partial G(\tilde{y},\tilde{y}_{1})}{\partial\tilde{y}}\frac{\partial G(\tilde{y},\tilde{y_{2}})}{\partial\tilde{y}}=\frac{\mathcal{G}(\tilde{y},\tilde{y}_{1},\tilde{y}_{2})}{\tilde{y}^{3-3w}}.$ (49)
Combining the above results, we could further simplify Eq.(III) as
$\mathcal{P}_{\mathrm{GW}}\left(y,qR_{*}\right)=\frac{\left[16\pi G\left(\bar{\tilde{e}}+\bar{\tilde{p}}\right)\bar{U}_{f}^{2}\right]^{2}(qR_{*})^{3}}{24\pi^{2}H^{2}H_{s}^{2}y^{4}}$
×$~{}~{}~{}~{}\times\int{\rm d}y_{-}\frac{\tilde{\Pi}^{2}\left(q|\eta_{1}-\eta_{2}|\right)}{\tilde{q}^{2}}\int{\rm d}y_{+}\frac{\mathcal{G}\left(\tilde{y},\tilde{y}_{1},\tilde{y}_{2}\right)}{y_{1}^{\frac{3+3w}{2}}y_{2}^{\frac{3+3w}{2}}}.$ (50)
It turns out that $y_{-}\ll y_{+}$ generally holds except in some special parameter spaces for sound shell model. This could be verified by noting that $\beta R_{*}=(8\pi)^{1/3}v_{w}\approx 3v_{w}$ and $R_{*}a_{s}H_{s}\sim\mathcal{O}(10^{-3})$, where $\beta$ is the value of the derivative of the logarithm of the probability of bubble nucleation with respect to time at some reference temperature Guo et al. (2021). Thus $y_{-}\sim\mathcal{O}(10^{-3})/v_{w}\times\beta(\eta_{1}-\eta_{2})$ according to Eq.(42). When $|\eta_{1}-\eta_{2}|$ exceeds $\mathcal{O}(10)$, the $\tilde{\Pi}$ quickly decays to zero and so the effective integral interval of $y_{-}$ is very close to zero Guo et al. (2021). On the contrary, $y_{+}\geq 1$ for the minimum of $y_{1}$ or $y_{2}$ is 1. Thus, we have $y_{+}\gg y_{-}$, which implies that we could keep the leading order of $y_{-}$ when dealing with the integration of $y_{+}$. The $\Upsilon$ factor is exactly the result of the integration of $y_{+}$ and it serves as a multiplicative factor in the expression of gravitational wave spectrum.
To calculate this complex integration, it is worth noting that the variable of the Green’s function Eq.(III) is the combination $\tilde{q}=q/H_{s}a_{s}=qR_{*}/H_{s}a_{s}R_{*}\sim\mathcal{O}(10^{3})qR_{*}$. This indicates that we could use the asymptotic expansion of the Bessel functions
$J_{\nu}(x)$ $\sim\frac{\sqrt{\frac{\pi}{2}}{{\rm cos}\left[x-\frac{1}{4}(1+2\nu)\right]}}{\sqrt{x}},$ (51)
if $\tilde{q}$ is large. Then the integrand takes the form of a fraction. After omitting the linear terms associated with $H_{s}a_{s}R_{*}$, the oscillation term of its numerator $\mathcal{N}_{\rm osc}$ can be expressed as
$\mathcal{N}_{\rm osc}=4\left[qR_{*}\right]^{2}y^{1+3w}(\tilde{P}_{1}\tilde{P}_{2}-\tilde{P}_{3}\tilde{P}_{4})(\tilde{P}_{5}\tilde{P}_{6}-\tilde{P}_{7}\tilde{P}_{8}),$ (52)
where $\tilde{P}_{i}$ are complex expressions involving trigonometric functions and they must be carefully handled to extract the correct results from strong oscillations. The key to significantly simplify these expressions lies in properly utilizing the fact that $c_{k}\equiv R_{*}a_{s}H_{s}\sim\mathcal{O}(10^{-3})$. After ignoring some small terms, we find that
$\tilde{P}_{1}$ $=\tilde{P}_{5}=AP_{1}=-2A{\rm cos}\left[\frac{3c_{k}\pi w-2qR_{*}y^{\frac{1+3w}{2}}}{c_{k}+3c_{k}w}\right],$ (53)
$\tilde{P}_{2}$ $=B{\rm cos}\left[\frac{\pi}{4}\left(1+\frac{6-6w}{2+6w}\right)-\frac{qR_{*}y_{+}}{c_{k}}+\frac{qR_{*}y_{-}}{2c_{k}}\right],$ (54)
$\tilde{P}_{3}$ $=\tilde{P}_{7}=AP_{3}=2A{\rm sin}\left[\frac{c_{k}\pi-2qR_{*}y^{\frac{1+3w}{2}}}{c_{k}+3c_{k}w}\right],$ (55)
$\tilde{P}_{4}$ $=B{\rm cos}\left[\frac{\pi}{4}\left(1+\frac{6w-6}{2+6w}\right)-\frac{qR_{*}y_{+}}{c_{k}}+\frac{qR_{*}y_{-}}{2c_{k}}\right],$ (56)
$\tilde{P}_{6}$ $=B{\rm cos}\left[\frac{\pi}{4}\left(1+\frac{6-6w}{2+6w}\right)-\frac{qR_{*}y_{+}}{c_{k}}-\frac{qR_{*}y_{-}}{2c_{k}}\right],$ (57)
$\tilde{P}_{8}$ $=B{\rm cos}\left[\frac{\pi}{4}\left(1+\frac{6w-6}{2+6w}\right)-\frac{qR_{*}y_{+}}{c_{k}}-\frac{qR_{*}y_{-}}{2c_{k}}\right],$ (58)
where
$A=\frac{\sqrt{c_{k}+3c_{k}w}}{\sqrt{\pi qR_{*}y^{\frac{1+3w}{2}}}},\quad B=\frac{\sqrt{2c_{k}}}{\sqrt{\pi qR_{*}y_{+}}}.$ (59)
By denoting
$x$ $=\frac{\pi}{4}\left(1+\frac{6-6w}{2+6w}\right)-\frac{qR_{*}y_{+}}{c_{k}},$ (60)
$x^{\prime}$ $=\frac{\pi}{4}\left(1+\frac{6w-6}{2+6w}\right)-\frac{qR_{*}y_{+}}{c_{k}},$ (61)
$z$ $=\frac{qR_{*}y_{-}}{2c_{k}},$ (62)
Eq.(52) can be written as
$\mathcal{N}_{\rm osc}$ $\sim A^{2}B^{2}\left[P_{1}*{\rm cos}(x+z)-P_{3}*{\rm cos}(x^{\prime}+z)\right]$
×$\times\left[P_{1}*{\rm cos}(x-z)-P_{3}*{\rm cos}(x^{\prime}-z)\right].$ (63)
Because we want to extract the term related to $y_{-}$, we use the properties of trigonometric functions
${\rm cos}(x+z){\rm cos}(x^{\prime}-z)=\frac{{\rm cos}(x+x^{\prime})+{\rm cos}(x-x^{\prime}+2z)}{2},$ (64)
to obtain the desired form of $\mathcal{N}_{\rm osc}$ which includes the sum of
$A^{2}B^{2}\left[P_{1}^{2}\frac{{\rm cos}(2x)}{2}-2P_{1}P_{3}\frac{{\rm cos}(x+x^{\prime})}{2}+P_{3}^{2}\frac{{\rm cos}(2x^{\prime})}{2}\right],$ (65)
and
$A^{2}B^{2}$ $\left[P_{1}^{2}\frac{{\rm cos}(2z)}{2}-P_{1}P_{3}{\rm cos}(2z){\rm cos}(x-x^{\prime})+P_{3}^{2}\frac{{\rm cos}(2z)}{2}\right].$ (66)
The denominator $\mathcal{D}$ is the form of
$\mathcal{D}=8c_{k}^{2}(M_{1}-M_{2}+M_{3}M_{4})(M_{5}-M_{6}+M_{7}M_{8}),$ (67)
where $M_{i}$ are also complex expressions involving trigonometric functions. By similar steps, the $\mathcal{D}$ could be simplified as
$\mathcal{D}=\frac{32c_{k}^{4}}{\pi^{2}}\frac{\left({\rm cos}\left[\frac{3w\pi-\pi}{3w+1}\right]-{\rm cos}\left[\frac{2\pi}{3w+1}\right]\right)^{2}}{(qR_{*}y^{+})^{2}}.$ (68)
Collecting all the above terms, the integral with respect to $y_{+}$ can ultimately be expressed as
$\int{\rm d}y_{+}\frac{\mathcal{G}}{y_{1}^{\frac{3+3w}{2}}y_{2}^{\frac{3+3w}{2}}}=\mathcal{G}_{1}+\mathcal{G}_{2},$ (69)
where
$\mathcal{G}_{1}$ $=\int{\rm d}y_{+}~{}\bigg{\{}\frac{\tilde{q}^{2}(1+3w)^{\frac{-4}{1+3w}}y_{+}^{\frac{-4}{1+3w}}2^{\frac{4}{1+3w}}}{({\rm cos}\left[\frac{3w\pi-\pi}{3w+1}\right]-{\rm cos}\left[\frac{2\pi}{3w+1}\right])^{2}}$
×$\times\left[P_{1}^{2}\frac{{\rm cos}(2x)}{2}+P_{3}^{2}\frac{{\rm cos}(2x^{\prime})}{2}-P_{1}P_{3}{\rm cos}\left(x+x^{\prime}\right)\right]\bigg{\}},$ (70)
and
$\mathcal{G}_{2}$ $=\int{\rm d}y_{+}~{}\bigg{\{}\frac{\tilde{q}^{2}(1+3w)^{\frac{-4}{1+3w}}y_{+}^{\frac{-4}{1+3w}}2^{\frac{4}{1+3w}}}{\left({\rm cos}\left[\frac{3w\pi-\pi}{3w+1}\right]-{\rm cos}\left[\frac{2\pi}{3w+1}\right]\right)^{2}}$
×$\times{\rm cos}\left(\frac{qR_{*}y_{-}}{c_{k}}\right)\left[\frac{P_{1}^{2}}{2}+\frac{P_{3}^{2}}{2}-P_{1}P_{3}{\rm cos}\left(x-x^{\prime}\right)\right]\bigg{\}}.$ (71)
Because of the cosine functions within the brackets of $\mathcal{G}_{1}$, the integrand curve of $\mathcal{G}_{1}$ oscillates rapidly, as displayed in Fig. (1) for example.
Figure 1: Comparing behaviors of the integrands of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. The blue solid line represents the integrand of $\mathcal{G}_{1}$, while the red solid line represents that of $\mathcal{G}_{2}$. In both plots, $qR_{*}$ is set to 10 and $c_{k}$ is set to 0.00079.
Thus, we expect its integral value just to provide the oscillatory behavior rather than the dominant term, and its integral value should be much smaller than the latter. This can be verified by using the Riemann-Lebesgue lemma if $y$ is large enough Stein and Murphy (1993). We also calculate its integral value over $\tilde{q}^{2}$ when $c_{k}=0.00079$ and $qR_{*}=10$ by numerical method and find that the corresponding value is 0.0286, -0.0035 and -0.0163 for $w=1/4$ and $y=5,10,50$, respectively, which is much smaller than the corresponding value of $\mathcal{G}_{2}$ in most intervals, see below.
Figure 2: The $\Upsilon$ factor for different $w$. In the plot, $c_{k}$ is set to 0.00079, and $qR_{*}$ is set to 10.
The $\mathcal{G}_{2}$ could be calculated directly in the limit of $y_{-}\to 0$ by noting that the term in the brackets is $y_{+}$ independent, and can be written in the following form:
$\mathcal{G}_{2}$ $\equiv\tilde{q}^{2}\frac{{\rm cos}(\tilde{q}y_{-})}{2}\Upsilon^{\nu\notin\mathbb{N}}(y;w),$ (72)
where
$\Upsilon^{\nu\notin\mathbb{N}}$ $=\frac{2\left[y^{3(w-1)/2}-1\right]}{3(w-1)\left({\rm cos}\left[\frac{3w\pi-\pi}{3w+1}\right]-{\rm cos}\left[\frac{2\pi}{3w+1}\right]\right)^{2}}$
×$\times\left(P_{1}^{2}+P_{3}^{2}-2P_{1}P_{3}{\rm cos}\left[\frac{3\pi(w-1)}{6w+2}\right]\right),$ (73)
and $\mathbb{N}$ is the set of integers. By applying trigonometric identities for further simplification, we can eliminate the trigonometric functions in both the numerator and the denominator, allowing us to obtain the following final expression for $\Upsilon$ when $\nu$ is not an integer,
$\Upsilon^{\nu\notin\mathbb{N}}=\frac{2\left[y^{3(w-1)/2}-1\right]}{3(w-1)}.$ (74)
For the case where $\nu$ is an integer, we can follow a similar simplification method to obtain the corresponding $\Upsilon$, which yields the same result as it in the non-integer case. Thus, our final expression for $\Upsilon$ is
$\Upsilon=\frac{2\left[y^{3(w-1)/2}-1\right]}{3(w-1)}.$ (75)
With this result, the final dimensionless gravitational wave spectrum is
$\mathcal{P}_{\rm GW}=\Upsilon(y;w)\frac{\left[16\pi G\left(\bar{\tilde{e}}+\bar{\tilde{p}}\right)\bar{U}_{f}^{2}\right]^{2}\left(qR_{*}\right)^{3}}{48\pi^{2}H^{2}H_{s}^{2}y^{4}}\int{\rm d}y_{-}{\rm cos}(\tilde{q}y_{-})\tilde{\Pi}^{2}.$ (76)
Compared with the result of Hindmarsh and Hijazi (2019), the $\Upsilon$ factor is an extra factor originated from the finite duration time of sound waves. The $\Upsilon$ factor for different values of $w$ is shown in Fig.(2). Obviously, the integral value of $\mathcal{G}_{1}$ over $\tilde{q}^{2}$ is indeed much smaller than the $\Upsilon$ factor, so it is safe to omit the contribution of $\mathcal{G}_{1}$. Considering the radiation-dominated universe case ($w=\frac{1}{3}$) and the matter-dominated universe case ($w=0$), our new $\Upsilon$ factor could correctly reproduce our previous results Guo et al. (2021)
$\Upsilon_{\rm RD}$ $=1-\frac{1}{y},$ (77)
$\Upsilon_{\rm MD}$ $=\frac{2}{3}\left(1-\frac{1}{y^{3/2}}\right).$ (78)
In our previous work, we referred to $\Upsilon$ as the suppression factor because its values are always lower than the corresponding asymptotic value. When extending to an arbitrary single-component universe, $\Upsilon$ can still act as a suppression factor for most of $w$, and the effect becomes even more pronounced when $w$ is sufficiently large, as shown by the blue line in Fig.(2). When the universe is dominated by the scalar field and some other matters, it is possible that the corresponding effective $w$ is about 2/3 and its $\Upsilon$ could be larger than one, even if the duration time is not very long. So it would indeed hold significant meaning if the source of gravitational waves went through some kinds of non-standard cosmological processes. In addition, this extension also makes strides in precisely calculating the gravitational wave spectrum during the transition period, when the primary constituents of the universe undergo changes.
An interesting aspect is that our derivation relies only on the conditions that the source is stationary and that the characteristic quantity $c_{k}$ is relatively small. Consequently, the resulting $\Upsilon$ factor is actually independent of the sound shell model. This suggests that $\Upsilon$ may serve as a more universal factor to describe how gravitational waves induced by a stationary source evolve over time Roper Pol et al. (2024).
It should be noted that our new $\Upsilon$ is ill-defined when $w=1$. In order to investigate the behavior of the $\Upsilon$ in this vicinity, we perform a Taylor expansion as below
$\left.\Upsilon(y;w)\right|_{w\to 1}$ $\approx{\rm ln}(y)+\mathcal{O}(w-1).$ (79)
We find that although $\Upsilon$ does not diverge in the vicinity of $w=1$, it behaves like a logarithmic function, suggesting that $\Upsilon$ tends to infinity when the active time of the source is sufficiently long. To understand the underlying reasons, we reexamine Eq. (III) and observe that the logarithmic behavior stems from the power of $y_{+}$ in the integrand, which equals $-1$ when $w=1$. This prompts us to analyze the composition of this power. There are three contributions to this power: (1) the Jacobian determinant provides a factor of $(y_{1}y_{2})^{\frac{1-3w}{2}}$; (2) the Green’s function contributes a factor of $y_{+}^{2}$; (3) the UETC of the energy-momentum tensor $\Pi^{2}$ introduces a factor of $y_{1}^{4}y_{2}^{4}$, which is canceled by $a^{2}(\eta_{1})a^{2}(\eta_{2})$, leaving a term $y_{1}^{2}y_{2}^{2}$, as shown in Eq.(II). Thus, we can express $\Upsilon$ roughly as
[…]
where we have assumed $y_{-}$ to be close to zero and so $y_{1}\approx y_{2}\approx y_{+}^{\frac{2}{3w+1}}$. In the above equation, the denominator represents the generation of gravitational waves induced by an instant source, the numerator reflects the source’s dilution effect, and the determinant accounts for the impact of the coordinate transformation. It is clear that the first two terms compete and cancel each other when $w=1$, leaving only the determinant to contribute a factor of $1/y_{+}$ and resulting in logarithmic behavior. This fact inspires us to revisit our coordinate system to obtain a qualitative explanation. For $w\neq-1$, scale factor $a$ is proportional to $t^{\frac{2}{1+3w}}$, and so
$y=\frac{a(t)}{a_{s}}=\left(\frac{t}{t_{s}}\right)^{\frac{2}{3(1+w)}}.$ (81)
If we fix the value of $y$, then larger values of $w$ will result in a longer duration for the slowly increasing behavior, as seen in Fig.(3). This is equivalent to the gravitational wave source dilutes more slowly, allowing more time for the production of gravitational waves, which causes the asymptotic value of $\Upsilon$ to increase with $w$, and ultimately results in logarithmic behavior when $w=1$.
Figure 3: Schematic diagram of $y$ as a function of time $t$.
One might ask how to deal with the infinity of $\Upsilon$ when $y\to+\infty$. In fact, we have implicitly assumed in the above procedure that when the gravitational wave source is active, the equation governing the background energy density remains unchanged. If the gravitational wave source lasts long enough, the dominant component of the cosmic background may change, which means that $w$ is actually a function of $y$. Therefore, we cannot simply take $y$ to infinity while fixing $w$ as a constant.
In addition, the universe is in the so-called kination period when $w=1$, during which the kinetic energy of a free scalar field dominates. Ref.Gouttenoire et al. (2021) demonstrated that this period must occur either immediately after inflation or following the matter-dominated era. Since the energy density during the kination period decreases rapidly as $a^{-6}$, the duration of this period is strictly restricted to be consistent with the results of Big Bang Nucleosynthesis and Cosmic Microwave Background. Thus, even if $w$ is always one, $y$ still cannot be taken as positive infinity.
Lastly, $\Upsilon$ also does not have an asymptotic value when $w>1$. However, due to the requirements of causality, this case can be disregarded. Specifically, the energy density will be less than the pressure density when $w>1$, leading to the speed of sound exceeds the speed of light.
IV GW spectrum today
With $\Upsilon$ obtained, the gravitational wave spectrum today can be written in the following factorized form:
$\Omega_{\text{GW}}h^{2}(f)|_{\text{today}}$ $=\text{Redshift factor}\times\mathcal{P}_{\rm GW},$ (82)
where $\mathcal{P}_{\rm GW}$ can be found in Eq.(76) and the redshift factor is
${\rm Redshift~{}factor}=\left(\frac{a_{e}}{a_{0}}\right)^{4}\left(\frac{H_{e}}{H_{0}}\right)^{2}h^{2},$ (83)
with $a_{e}$ and $H_{e}$ representing the scale factor and Hubble parameter at the moment when gravitational waves cease to be produced, $a_{0}$ and $H_{0}$ denoting the current scale factor and Hubble parameter, while $h\approx 0.678$ being the Hubble parameter today in unit of $100{\rm km}/s/{\rm Mpc}$.
Expanding Eq.(76) and using the explicit expressions of $H$ and $H_{s}$, we can rewrite it as
$\mathcal{P}_{\rm GW}=3\Gamma^{2}\chi^{2}\bar{U}_{f}^{4}H_{s}a_{s}R_{*}\mathcal{A}\mathcal{S_{\rm SW}}\frac{y^{3(w+1)}}{y^{4}}\Upsilon(y;w),$ (84)
where $\chi$ represents the ratio of the source energy density to the total energy density of the universe at the source active time $t_{s}$, $\Gamma$ is the ratio of source enthalpy density to the corresponding energy density and $A\mathcal{S_{\rm SW}}$ is the spectrum function that denotes the integration
$\mathcal{A}\mathcal{S_{\rm SW}}=\frac{\left(qR_{*}\right)^{3}}{4\pi^{2}}\int{\rm d}y_{-}{\rm cos}(\tilde{q}y_{-})\tilde{\Pi}^{2}(y_{-},qR_{*}).$ (85)
In principle, the spectrum function can be derived by either calculating the integration described above or utilizing the results from the numerical simulations. For radiation-dominated era, the numerical result is
$\mathcal{A}\mathcal{S_{\rm SW}}=0.008\times\left(\frac{f}{f_{\rm SW}}\right)^{3}\left[\frac{7}{4+3\left(f/f_{\rm SW}\right)^{2}}\right]^{7/2},$ (86)
where $f_{\rm SW}$ is the peak frequency of the corresponding gravitational wave Hindmarsh et al. (2017).
In phase transition, we can further express $R_{*}$ by the inverse transition duration time $\beta$ and so the characteristic quantity $H_{s}a_{s}R_{*}$ can be rearranged as
$a_{s}H_{s}R_{*}\approx\left(8\pi\right)^{1/3}v_{w}\frac{H_{s}}{\beta},$ (87)
where $v_{w}$ is the bubble wall velocity and we have assumed that bubbles disappear quickly. With all the above ingredients, the final spectrum is
$\Omega_{\text{GW}}h^{2}(f)|_{\text{today}}$ $=4.04\Gamma^{2}\chi^{2}\bar{U}_{f}^{4}v_{w}\frac{H_{s}}{\beta}\mathcal{A}\mathcal{S}_{\rm SW}$
×$\times\left(\frac{a_{e}}{a_{0}}\right)^{4}\left(\frac{H_{e}}{H_{0}}\right)^{2}\frac{y^{3(w+1)}}{y^{4}}\Upsilon(y;w).$ (88)
It is interesting to note the explicit dependence of the gravitational wave spectrum on $w$. Finally, we emphasize that if the duration of gravitational wave production is sufficiently long that the dominant component of the cosmic background changes, the above spectrum is no longer valid and the corresponding $\Upsilon$ must be applied for each era accordingly.
V conclusion
In this work, we successfully utilized the sound shell model in an expanding universe to derive the factor $\Upsilon$, which reflects the growth of gravitational waves induced by sound waves in a single-component universe that satisfies the strong energy condition. We obtained the following observations: (1) Our new $\Upsilon$ accurately reproduces previous results for radiation-dominated and matter-dominated universes; (2) During the kination period, the new $\Upsilon$ exhibits a logarithmic behavior, driven by the competition between the rate of the gravitational wave production and the dilution rate of the gravitational wave source; (3) The new $\Upsilon$ does not depend on the sound shell model, suggesting that it may represent a more universal result and is not limited to cosmological phase transitions.
Acknowledgments
We would like to thank Ligong Bian and Rui-Xiao Zhang for helpful discussions. This work is supported by the National Natural Science Foundation of China (12105248, 11821505, 12075300, 12335005, 12147103), by Peng-Huan-Wu Theoretical Physics Innovation Center (12047503), and by the Research Fund for Outstanding Talents from Henan Normal University (5101029470335).